# SA-Person: Text-Based Person Retrieval with Scene-aware Re-ranking

### v2

@@ Abstract
Text-based person retrieval aims to identify a target individual from a gallery of images based on a natural language description. It presents a significant challenge due to the complexity of real-world scenes and the ambiguity of appearance-related descriptions. Existing methods primarily emphasize appearance-based cross-modal retrieval, often neglecting the contextual information embedded within the scene, which can offer valuable complementary insights for retrieval. To address this, we introduce ScenePerson-13W, a large-scale dataset featuring over 100,000 scenes with rich annotations covering both pedestrian appearance and environmental cues. Based on this, we propose SA-Person, a two-stage retrieval framework. In the first stage, it performs discriminative appearance grounding by aligning textual cues with pedestrian-specific regions. In the second stage, it introduces SceneRanker, a training-free, scene-aware re-ranking method leveraging multimodal large language models to jointly reason over pedestrian appearance and the global scene context. Experiments on ScenePerson-13W validate the effectiveness of our framework in challenging scene-level retrieval scenarios. The code and dataset will be made publicly available.
00footnotetext: * These authors contributed equally to this work.00footnotetext: †Corresponding author.
1 Introduction
@@
Person retrieval galiyawala2021person aims to identify a target individual within the image galleries based on a query, playing a vital role in applications like video surveillance and human-centric artificial intelligent systems li2018richly ; galiyawala2021person ; li2023dcel . Early efforts focus on image-based paradigms zhang2016learning ; liu2018pose ; sarfraz2018pose ; xiao2017joint ; he2018end ; yan2021anchor ; cao2022pstr ; yan2022exploring using cropped images as queries, and then shift toward text-based retrieval li2017person ; yu2019cross ; liu2023survey due to the accessibility of textual descriptions. Advances in vision-language pretraining (VLP) models, such as CLIP radford2021learning , also accelerate its development yang2023towards ; li2024adaptive ; zuo2024ufinebench ; luo2025graph ; liu2025dm ; jiang2023cross ; bai2023rasa ; cao2024empirical ; ergasti2024mars ; qin2024noisy . However, text-based methods rely solely on cropped images and textual descriptions of appearance, limiting their effectiveness in real-world scenarios. Recent methods zhang2023text ; su2024maca ; yan2025fusionsegreid extend retrieval to full-scene images but still focus narrowly on appearance, neglecting critical environmental scene cues. This limitation poses challenges in complex scenarios with similar-looking individuals, or the same person under varying contexts. Incorporating contextual scene semantics alongside individual appearance is essential for accurate and robust person retrieval.
@@
This work focuses on scene-aware text-based person retrieval. The first challenge is the lack of a dedicated dataset for this task. To this end, we introduce ScenePerson-13W, a novel dataset designed for real-world scene-aware person retrieval. It includes over $131,438$ images from $12,562$ YouTube video sequences, featuring diverse individuals across varying crowd densities, camera angles, and environments. Unlike existing datasets that focus on cropped images or weakly grounded captions, ScenePerson-13W provides rich image-text pairs with detailed descriptions encompassing both visual appearance and contextual scene information, which promotes text-to-person alignment.
@@
Building on ScenePerson-13W, we propose SA-Person, a two-stage framework for scene-aware person retrieval. In the first stage, appearance-related information is extracted from textual descriptions to identify candidate pedestrians via cross-modal matching, reducing the retrieval space with discriminative appearance features. In the second stage, we introduce SceneRanker, a training-free, scene-aware re-ranking module powered by MLLM. By employing structured multimodal prompts, SceneRanker enables holistic reasoning over appearance and global scene context, allowing effective disambiguation of visually similar individuals and significantly improving retrieval performance.
@@
  • We introduce ScenePerson-13W, a new large-scale dataset for scene-aware text-based person retrieval, with compositional descriptions that include both pedestrian appearance and global scene context, alleviating the limitations of prior datasets.
@@
  • We propose SceneRanker, the first re-ranking strategy for text-based person retrieval powered by MLLM, which follows a training-free pattern and incorporates visual grounding and contextual reasoning to improve retrieval accuracy in complex scenes.
@@
  • Extensive experiments on ScenePerson-13W demonstrate the effectiveness and generalizability of our proposed SA-Person across various retrieval settings and model backbones.
@@
Figure 1: Overview of the ScenePerson-13W construction pipeline. The construction pipeline involves scene segmentation, pedestrian detection and tracking, completeness filtering, image deduplication, and textual description generation. For each retained pedestrian, a description is generated based on the full image with a highlighted target, capturing their appearance, spatial location, and relationships with surrounding elements.
@@
Table 1: Statistics of person retrieval datasets. SCENEPERSON-13W shows a substantial advancement over existing person retrieval datasets by combining fine-grained appearance details with rich global scene contexts.
@@
Market-1501 zheng2015scalable Image (Cropped) Image (Cropped) 32,668
CUHK03 li2014deepreid Image (Cropped) Image (Cropped) 14,097
CUHK-SYSU xiao2017joint Image (Cropped) Image (Full) 18,184
PRW zheng2017person Image (Cropped) Image (Full) 11,816
CUHK-PEDES li2017person Text (Appearance) Image (Cropped) 80,412 40,206
ICFG-PEDES ding2021semantically Text (Appearance) Image (Cropped) 54,522 54,522
CUHK-SYSU-TBPS zhang2023text Text (Appearance) Image (Full) 35,960 18,184
PRW-TBPS zhang2023text Text (Appearance) Image (Full) 19,009 11,246
@@ 2 ScenePerson-13W Dataset
The ScenePerson-13W dataset is proposed to support scene-aware person retrieval by incorporating rich scene context alongside detailed pedestrian appearance in textual descriptions. It contains $131,438$ pedestrian instances across $112,872$ diverse scenes, with each instance accompanied by a textual description, a full-scene image, and a bounding box highlighting the pedestrian’s location. The dataset is split into a training set of $121,438$ images and a testing set of $10,000$ images, which preserves the complex scene structures and identity distributions while reducing redundancy and ensures comprehensive evaluation across diverse scenarios. The construction process of the dataset illustrated in Figure 1, involves five key steps: scene segmentation, pedestrian detection and tracking, completeness filtering, deduplication in image collection, and textual description generation. More examples from the dataset are presented in Appendix A.3.
@@
Image Collection. The image data in ScenePerson-13W is constructed from the QVHighlights video dataset lei2107qvhighlights through a multi-stage processing pipeline designed to ensure data quality and diversity. First, videos are segmented into distinct scenes, each representing a unique environmental context. This segmentation enables the dataset to capture a wide range of pedestrian-environment combinations, supporting comprehensive scene-aware analysis. Second, the pedestrian detection and tracking are performed using YOLOv11 khanam2024yolov11 . For each detected pedestrian instance, the frame with the highest confidence score is selected, and the corresponding full scene image with cropped pedestrian region are retained. Third, to enhance data quality, incomplete or noisy detections–such as those capturing only partial body parts–are filtered out using MediaPipe keypoint detection mediapipe_pose . Only images containing a head and at least one shoulder are preserved, ensuring the availability of essential visual features. Finally, a deduplication process is applied to reduce redundancy. Using CLIP radford2021learning , visual similarity between images is assessed. Instances that are highly similar in both the cropped pedestrian region and the full scene are discarded, promoting the uniqueness and diversity of the resulting dataset. Finally, we collect $131,438$ pedestrian instance, accompanied with $131,438$ corresponding scene images.
@@
Description Generation. For each of the $131,438$ retained pedestrian instances, a corresponding textual description is generated using InternVL2.5-8B chen2024expanding . Specifically, the model is prompted with the full-scene image in which the target pedestrian is highlighted using a red bounding box. The prompt is structured as follows:
@@
Generate a description of the individual within the red bounding box and the connection with the surroundings, in a style similar to the template: {template}. Ensure that the description is concise within $50$ words and accurate. Do not include text details or labels, and strictly avoid any punctuation other than commas and periods.
@@
Compared to existing person retrieval datasets, ScenePerson-13W offers a significant advancement by integrating fine-grained appearance details with rich global scene context across diverse scenarios, as shown in Table  1.
@@
3 Method
To tackle the challenges of holistic understanding in both pedestrian appearance and scene context, we propose SA-Person, a two-stage framework, as shown in Figure 2. The first stage aligns pedestrian-specific regions with appearance-related descriptions to retrieve candidates. The second stage refines the initial candidates by incorporating global scene context.
3.1 Appearance-based Retrieval
Appearance Embedding Training. The first stage focuses on retrieving a shortlist of candidate individuals from the gallery set based solely on their visual appearance, given the appearance-related content of the textual query. For a gallery of full-scene images $\mathcal{G}=\{I_{1},I_{2},\dots,I_{N}\}$ and a textual query $T$ describing the target person, we begin by extracting the appearance-related content from $T$. Specifically, we parse each query description in the dataset and extract the segments of the description that describes the target person’s visual traits, such as clothing, hairstyle, or accessories (e.g., “A man wearing a red jacket and blue jeans is laughing”), denoted as ${T}_{\text{app}}$. This ensures that the retrieval process prioritizes appearance-focused matching, excluding contextual or environmental cues at this stage.

### v3

@@ Abstract
Text-based person retrieval aims to identify a target individual from an image gallery using a natural language description. Existing methods primarily focus on appearance-driven cross-modal retrieval, yet face significant challenges due to the visual complexity of scenes and the inherent ambiguity of textual descriptions. The contextual information, such as landmarks and relational cues, provides complementary cues that can offer valuable complementary insights for retrieval, but remains underexploited in current approaches. Motivated by this limitation, we propose a novel paradigm: scene-aware text-based person retrieval, which explicitly integrates both individual appearance and global scene context to improve retrieval accuracy. To support this, we first introduce ScenePerson-13W, a large-scale benchmark dataset comprising over 100,000 real-world scenes with rich annotations encompassing both pedestrian attributes and scene context. Based on this dataset, we further present SA-Person, a two-stage retrieval framework. In the first stage, SA-Person performs discriminative appearance grounding by aligning textual descriptions with pedestrian-specific regions. In the second stage, it introduces SceneRanker, a training-free, scene-aware re-ranking module that refines retrieval results by jointly reasoning over pedestrian appearance and the global scene context. Extensive experiments on ScenePerson-13W and existing benchmarks demonstrate the effectiveness of our proposed SA-Person. Both the dataset and code will be publicly released to facilitate future research.
I Introduction
@@
Person retrieval [galiyawala2021person] aims to identify a target individual within the image galleries based on a query, playing a vital role in applications like video surveillance and human-centric artificial intelligent systems [li2018richly, galiyawala2021person, li2023dcel]. Early efforts focus on image-based paradigms [zhang2016learning, liu2018pose, sarfraz2018pose, xiao2017joint, he2018end, yan2021anchor, cao2022pstr, yan2022exploring] using cropped images as queries, and then shift toward text-based retrieval [li2017person, yu2019cross, liu2023survey] due to the accessibility of textual descriptions. Advances in large-scale Vision-Language Pretraining (VLP) models, such as CLIP [radford2021learning], also accelerate the development of text-based person retrieval [yang2023towards, li2024adaptive, zuo2024ufinebench, luo2025graph, liu2025dm, jiang2023cross, bai2023rasa, cao2024empirical, ergasti2024mars, qin2024noisy]. However, methods for text-based person retrieval rely solely on cropped images and appearance-focused textual descriptions, limiting their effectiveness in real-world scenarios. Recent methods [zhang2023text, su2024maca, yan2025fusionsegreid] extend retrieval to full-scene images but still focus narrowly on appearance, neglecting critical scene cues.
This limitation poses challenges in complex scenarios with similar-looking individuals, or the same person under varying contexts, as illustrated in Figure 1(a). In such cases, appearance-only retrieval often produces ambiguous or incorrect matches. By contrast, human descriptions about a target person are rarely restricted to appearance; instead, they naturally incorporate auxiliary contextual cues that complement visual traits, such as surrounding environments, landmark buildings, and co-occurring entities. As illustrated in Figure 1(b), incorporating scene context substantially enhance disambiguation beyond what appearance features alone can offer. To quantify the potential of scene context, we perform a preliminary analysis on the CUHK-SYSU dataset augmented with MLLM-generated descriptions. As shown in Figure 1(c), MLLMs with context-enriched text can achieve retrieval performance beyond the limits of fine-tuned domain-specific models. A comprehensive discussion is provided in Section III. Motivated by this observation, we focus on scene-aware text-based person retrieval, which explicitly incorporates both individual appearance and global scene context for more accurate and robust retrieval.
@@
The first challenge is the lack of a dedicated dataset for this task. To this end, we introduce ScenePerson-13W, a novel dataset designed for real-world scene-aware person retrieval. It includes over $131,438$ images from $12,562$ YouTube video sequences, featuring diverse individuals across varying crowd densities, camera angles, and environments. As summarized in Table II, existing datasets[zheng2015scalable, li2014deepreid, xiao2017joint, zheng2017person, li2017person, ding2021semantically, zhang2023text] typically focus on cropped images or weakly grounded captions, ScenePerson-13W provides rich image-text pairs with detailed descriptions encompassing both visual appearance and scene context, which promotes text-to-person alignment.
@@
Building on ScenePerson-13W, we propose SA-Person, a two-stage framework for scene-aware person retrieval. In the first stage, appearance-related information is extracted from textual descriptions to identify candidate pedestrians via cross-modal matching, reducing the retrieval space with discriminative appearance features. In the second stage, we introduce SceneRanker, a training-free, scene-aware re-ranking module that refines retrieval results by jointly reasoning over pedestrian appearance and global scene context, enabling holistic understanding to effectively disambiguate visually similar individuals and significantly improve retrieval performance.
Figure 1: Overview of person retrieval challenges and scene-aware insights. (a) Illustration of appearance-based retrieval limitations, where multiple candidates match the text A girl wearing a green and white striped shirt is smiling due to similar appearances. (b) Demonstration of scene-aware reranking, ranking candidate images by aligning scene context like located in a restaurant, near a window with red text with the text. (c) Retrieval performance across three progressive input configurations on the augmented CUHK-SYSU dataset. The three configurations are: (1) Cropped + App (cropped image with appearance-only text), (2) Full + App (full image with appearance-only text), and (3) Full + Context (full image with context-enriched text).
@@
  1. 1)
    We introduce ScenePerson-13W, the first dataset enabling a new retrieval paradigm of scene-aware text-based person retrieval by providing compositional descriptions that include both pedestrian appearance and global scene context, alleviating the limitations of prior datasets.
  2. 2)
@@
    We propose SceneRanker, the first re-ranking strategy for text-based person retrieval that can be seamlessly integrated as a plug-and-play module, incorporating visual grounding and contextual reasoning to disambiguate and enhanace retrieval performance in complex scenes.
  4. 4)
@@
    Extensive experiments on ScenePerson-13W and diverse benchmarks demonstrate the effectiveness and generalizability of our proposed SA-Person across various retrieval settings and model backbones.
Figure 2: Qualitative visualization of retrieval results across three progressive input configurations on the augmented CUHK-SYSU dataset. The blue frame contains the top candidates from the domain-specific model IRRA (ViT-L/14), and the orange frame contains the top candidates from the MLLM (InternVL-8B). The green-bordered image is the ground truth.
II Related works
II-A Text-Based Person Retrieval
Text-Based Person Retrieval (TBPR) aims to retrieve specific individuals based on natural language descriptions, typically from a gallery of cropped person images [yang2023towards, li2024adaptive, zuo2024ufinebench, luo2025graph, liu2025dm]. These methods focus on fine-grained cross-modal alignment between visual appearances and textual attributes, such as clothing, gender, or accessories [shu2022see, zuo2024ufinebench, wang2024fine, wei2024fine, zhao2025cross]. With the advent of vision-language models like CLIP [radford2021learning], TBPR performance has significantly improved. For instance, IRRA [jiang2023cross] introduces a multi-modal interaction encoder for implicit relational reasoning, and TBPS-CLIP [cao2024empirical] proposed a simple yet strong baseline through enhanced training strategies. To address noise in cross-modal data, RaSa [bai2023rasa] proposed sensitivity-aware and relation-aware representation learning for robustness against weakly matched pairs; MARS [ergasti2024mars] incorporated visual reconstruction and attribute-sensitive losses to mitigate inter-identity noise and intra-identity variations; RDE [qin2024noisy] introduced dual embeddings to handle under-correlated or mislabeled image-text pairs under noisy conditions. Beyond cropped-image retrieval, some methods [zhang2023text, su2024maca, yan2025fusionsegreid] retrieval individuals directly from unprocessed full-scene images, where individuals must be located and identified in complex visual environments Specifically, Zhang et al. [zhang2023text] integrated textual cues into region proposals, enabling text descriptions to guide the detection process. Su et al. [su2024maca] introduced a memory-aided coarse-to-fine alignment strategy, using memory banks to enhance cross-modal contrastive learning. Yan et al. [yan2025fusionsegreid] explored multimodal fusion by combining textual and visual features with precise segmentation techniques to boost retrieval performance. While these methods often overlook the rich contextual information present in the surrounding scene, which could offer valuable clues for more accurate and robust person retrieval.
II-B Multimodal Large Language Models for Scene and Region-Level Reasoning
Recent advances in Multimodal Large Language Models (MLLMs) enabled powerful visual reasoning across complex scenes. Current efforts primarily progresses along two directions: comprehensive scene understanding using global image-text inputs [chen2024sharegpt4v, chen2024expanding, bai2025qwen2, touvron2023llama, mitra2024compositional, fan2024mllm], and fine-grained region-level grounding [guo2024regiongpt, zhou2023regionblip, peng2023kosmos, chen2023shikra, ma2024groma, rasheed2024glamm]. For comprehensive scene understanding, recent works [mitra2024compositional, fan2024mllm] enables MLLMs to explicitly represent structured and compositional semantics. For instance, Mitra et al.  [mitra2024compositional] proposed generating scene graphs via MLLMs and incorporating them into prompts, facilitating compositional reasoning without relying on annotated scene graphs; MLLM-SUL  [fan2024mllm] introduced a dual-branch encoder and a fine-tuned LLaMA [touvron2023llama] model to jointly perform scene description in traffic scenarios. To support region-level grounding, recent studies  [guo2024regiongpt, zhou2023regionblip, chen2023shikra] enhance MLLMs with explicit spatial conditioning. For example, RegionGPT [guo2024regiongpt] and RegionBLIP [zhou2023regionblip] have been proposed adopt region-wise feature pooling within bounding boxes or masks to extract localized representations for improved spatial reasoning; KOSMOS-2 [peng2023kosmos] and Shikra [chen2023shikra] encode spatial positions as discrete tokens or numeric coordinates to support referential understanding; GLaMM [rasheed2024glamm] enables grounded interaction through joint generation of text and segmentation masks.
In this work, we combine the complementary strengths of domain-specific TBPR models and MLLMs for scene-aware text-based person retrieval, fusing efficient appearance alignment with holistic context reasoning to robustly disambiguate in complex scenes.
TABLE I: Retrieval performance across three progressive input configurations on CUHK-SYSU dataset, augmented using MLLM-generated descriptions.
Cropped Image Full Image Full Image
& Appearance-only Text & Appearance-only Text & Context-enriched Text
Model R@1 R@5 R@10 mAP R@1 R@5 R@10 mAP R@1 R@5 R@10 mAP
Domain-specific TBPR models (zero-shot)
RDE [qin2024noisy] (ViT-B/16) 40.62 79.19 84.18 57.34 21.17 50.09 62.28 35.19 31.82 69.58 81.32 48.53
IRRA [jiang2023cross] (ViT-B/16) 41.62 79.45 84.73 58.27 23.23 53.47 66.09 37.76 35.78 74.68 85.39 53.04
Domain-specific TBPR models (fine-tuned)
IRRA [jiang2023cross] (ViT-B/16) 47.60 84.00 89.14 63.63 33.76 73.69 84.40 51.37 66.61 98.53 99.60 80.72
RDE [qin2024noisy] (ViT-B/16) 48.73 83.78 88.88 63.93 35.12 73.47 82.97 51.98 67.23 98.31 99.60 81.05
IRRA [jiang2023cross] (ViT-L/14) 50.57 85.54 89.36 66.01 36.81 76.66 86.13 54.63 71.41 98.86 99.93 83.56
Multimodal large model (zero-shot)
InternVL-8B [chen2024expanding] 23.38 40.92 50.79 31.03 46.32 80.37 92.13 59.39 85.27 97.53 98.90 90.63
III Preliminary Analysis of Current Models on Text-Based Person Retrieval
Although text-based person retrieval has made significant strides driven by vision-language pretraining models, existing methods [jiang2023cross, bai2023rasa, cao2024empirical, ergasti2024mars, qin2024noisy] predominantly rely on appearance descriptions and cropped pedestrian images, overlooking the potential value of scene context. To empirically substantiate this oversight and demonstrate the complementary strengths of domain-specific TBPR models and MLLMs in handling increasing scene complexity, which thereby motivating our two-stage SA-Person framework, we conduct a preliminary analysis on the augmented CUHK-SYSU [xiao2017joint].
III-A Analysis Configuration
Originally featuring full-scene images with annotated target pedestrian bounding boxes, we augment CUHK-SYSU with MLLM-generated textual descriptions, with two types per target: (1) appearance-only descriptions, capturing visual attributes such as clothing, accessories, and pose; and (2) context-enriched descriptions, augmenting these with contextual details that include surroundings, spatial relations and interactions. This enables a gradual incorporation of scene context, tracing the shift from appearance-focused to context-aided retrieval through three progressive input configurations:
  1. 1)
    Appearance-only text & cropped image: Relies solely on appearance-focused descriptions and the target’s cropped bounding box region, isolating appearance alignment.
  2. 2)
    Appearance-only text & full image: Pairs the same descriptions with complete scene images, introducing scene information without explicit contextual guidance.
  3. 3)
    Context-enriched text & full image: Combines augmented descriptions with full scenes, evaluating joint reasoning over appearance and environmental semantics.
Evaluated models span two architectural paradigms: (1) state-of-the-art domain-specific TBPR models, including IRRA [jiang2023cross] and RDE [qin2024noisy], which use dual-encoder frameworks. For this analysis, we finetune these models on the augmented CUHK-SYSU dataset, adapting to each configuration’s input modality (e.g., holistic visuals for the last setting). We also evaluate their zero-shot performance using pre-trained weights from CUHK-PEDES. (2) Multimodal large language models (MLLMs), such as InternVL-8B [chen2024expanding], which support zero-shot retrieval without modality-specific adaptation. However, current MLLMs struggle with gallery-wide retrieval across vast image galleries, constrained by their bounded context capacities that restrict simultaneous handling of extensive images [niu2025chatreid]. To address this, we transform the complex retrieval task into a multi-round scoring and ranking process through engineering enhancements.
III-B Performance Analysis
Table I presents results across the three progressive configurations, revealing critical insights into the role of scene context in TBPR and highlighting the complementary strengths of domain-specific TBPR models and MLLMs when handling scene complexity.
In the appearance-only text & cropped image setting, domain-specific models dominate. Zero-shot IRRA pre-trained on CUHK-PEDES achieves 41.62% R@1, rising to 50.57% after fine-tuning, both surpassing zero-shot InternVL-8B’s 23.38% R@1. Their specialized architecture and extensive appearance-focused pretraining yield highly discriminative embeddings for clean, isolated regions. In contrast, zero-shot MLLMs lack domain alignment, producing coarser cross-modal representations.
However, switching to appearance-only text & full image setting, full scenes without contextual support, domain-specific models exposes a sharp vulnerability. IRRA with ViT-L/14 drops from 50.57% to 36.81% R@1, and zero-shot domain models fall to $\sim$23%. Distracting environmental elements like cluttered backgrounds and occlusions overwhelm appearance-based embeddings lacking semantic safeguards. In contrast, zero-shot InternVL-8B rises to 46.32% R@1, leveraging inherent scene understanding to mitigate interference.
Shifting to the context-enriched text & full image setting further underscores the value of scene context, where contextual augmentation markedly enhances retrieval accuracy across models. As shown in Figure 2, the contextual support enables refined disambiguation. Zero-shot IRRA yields only 35.78% R@1 due to limited generalization to context-enriched descriptions. Fine-tuning boosts R@1 to 71.41%, a 99% relative gain, yet still falls short of the zero-shot MLLM baseline, where InternVL-8B attains 85.27% R@1, demonstrating clear superior inherent generalization and seamless integration of local appearance with global elements like landmarks and relational cues. Domain-specific models, despite their improvements, cannot fully leverage these aspects.
III-C Time Complexity Analysis
For computational cost, MLLMs incur substantially higher expenses than domain-specific models due to the multi-round gallery retrieval process. Let $\mu_{s}$ and $\mu_{m}$ denote the per-image time units for domain-specific models and MLLMs, respectively, with $\mu_{m}\approx 2$ orders of magnitude larger than $\mu_{s}$ due to intensive multimodal processing. Domain-specific models require $O(N\mu_{s})$ for a gallery of size $N$. MLLMs demand an initial $O(N)$ traversal for match scores at $O(N\mu_{m})$ time, plus $O(M/B)$ batch ranking operations with per-batch time $T_{b}=\mu_{m}\times B$, where $M$ is the candidate set size and $B$ is the batch size. The total time thus simplifies to $O(N\mu_{m}+K\mu_{m})$. This yields an overall per-query time orders of magnitude higher than that of domain-specific models, scaling linearly with $N$ and limiting scalability for large galleries.
In summary, this analysis reveals a key contrast: domain-specific models excel in straightforward appearance tasks but demonstrate poor zero-shot generalization and struggle with scene variations, while MLLMs deliver superior and robust holistic reasoning in context-rich environments at a higher computational cost. The widening performance gaps as scene context increases underscore the value of approaches that connect detailed region matching and broader reasoning, directly motivating our development of SCENEPERSON-13W and the two-stage SA-Person framework to harness these complementary strengths for robust, real-world TBPR.
@@ III-C Time Complexity Analysis
Figure 3: Overview of the ScenePerson-13W construction pipeline. The construction pipeline involves scene segmentation, pedestrian detection and tracking, completeness filtering, image deduplication, and description generation. For each retained pedestrian, a description is generated based on the full image with a highlighted target, capturing their appearance, spatial location, and relationships with surrounding elements.
@@
TABLE II: Statistics of person retrieval datasets. SCENEPERSON-13W shows a substantial advancement over existing person retrieval datasets by combining fine-grained appearance details with rich global scene contexts.
@@
  Market-1501 [zheng2015scalable]   Image (Cropped)   Image (Cropped)   –   32,668
  CUHK03 [li2014deepreid]   Image (Cropped)   Image (Cropped)   –   14,097
  CUHK-SYSU [xiao2017joint]   Image (Cropped)   Image (Full)   –   18,184
  PRW [zheng2017person]   Image (Cropped)   Image (Full)   –   11,816
  CUHK-PEDES [li2017person]   Text (Appearance)   Image (Cropped)   80,412   40,206
  ICFG-PEDES [ding2021semantically]   Text (Appearance)   Image (Cropped)   54,522   54,522
  CUHK-SYSU-TBPS [zhang2023text]   Text (Appearance)   Image (Full)   35,960   18,184
  PRW-TBPS [zhang2023text]   Text (Appearance)   Image (Full)   19,009   11,246
@@ IV ScenePerson-13W Dataset
The ScenePerson-13W dataset is proposed to support scene-aware text-based person retrieval by incorporating rich scene context alongside detailed pedestrian appearance in textual descriptions. It contains $131,438$ pedestrian instances across $112,872$ diverse scenes, with each instance accompanied by a textual description, a full-scene image, and a bounding box highlighting the pedestrian’s location. The dataset consists of $121,438$ training images and $10,000$ test images, which preserves scene complexity and pedestrian distributions with reduced redundancy, enabling comprehensive evaluation across diverse scenarios.
The construction process of the dataset illustrated in Figure 3, involves five key steps: scene segmentation, pedestrian detection and tracking, completeness filtering, deduplication in image collection, and textual description generation.
@@
Image Collection. The image data in ScenePerson-13W is constructed from the QVHighlights video dataset [lei2107qvhighlights] through a multi-stage processing pipeline designed to ensure data quality and diversity. 1) Scene segmentation. Videos are segmented into distinct scenes, each representing a unique environmental context. This segmentation enables the dataset to capture a wide range of pedestrian-environment combinations, supporting comprehensive scene-aware analysis. 2) Pedestrian detection and tracking. Each segmented scene is then processed to detect and track pedestrians. For each detected pedestrian instance, the frame with the highest confidence score is selected, and the corresponding full scene image with cropped pedestrian region are retained. 3) Completeness filtering. To enhance data quality, incomplete or noisy detections–such as those capturing only partial body parts–are filtered out using MediaPipe keypoint detection [mediapipe_pose]. Only images containing a head and at least one shoulder are preserved, ensuring the availability of essential visual features. 4) Deduplication. Finally, a deduplication process is applied to reduce redundancy. Using CLIP [radford2021learning], visual similarity between images is assessed. Instances that are highly similar in both the cropped pedestrian region and the full scene are discarded, promoting the uniqueness and diversity of the resulting dataset. Finally, we collect $131,438$ pedestrian instance, accompanied with $131,438$ corresponding scene images.
Figure 4: Sunburst visualization of ScenePerson-13W, showing hierarchical feature distribution across Appearance, Landmarks, and Proximity.
@@
Description Generation. For each of the $131,438$ retained pedestrian instances, a corresponding textual description is generated using InternVL2.5-8B [chen2024expanding]. Specifically, the model is prompted with the full-scene image in which the target pedestrian is highlighted using a red bounding box. The prompt is structured as follows:
@@
Generate a description of the individual within the red bounding box and the connection with the surroundings. Ensure that the description is a natural, fluent paragraph (under $50$ words) that a real witness would actually say to police or friends when trying to help find this exact person. Do not include text details or labels, and strictly avoid any punctuation other than commas and periods.
@@
Compared to existing person retrieval datasets[zheng2015scalable, li2014deepreid, xiao2017joint, zheng2017person, li2017person, ding2021semantically, zhang2023text], ScenePerson-13W offers a significant advancement by integrating fine-grained appearance details with rich global scene context across diverse scenarios and surpassing prior datasets in scale, as shown in Table  II. And figure 4 further illustrates the hierarchical feature distribution in ScenePerson-13W, enabling precise, context-rich descriptions that jointly capture fine-grained appearance and complex scene context in real-world settings.
@@
V Method
Motivated by the insights revealed in Section III, we propose SA-Person, a straightforward yet effective two-stage framework, as shown in Figure 5. The first stage aligns pedestrian-specific regions with appearance-related textual descriptions to retrieve a compact set of candidates. The second stage refines the top candidates by incorporating global scene context via a training-free scene-aware re-ranking method.
V-A Appearance-based Retrieval
Appearance Embedding Training. The first stage focuses on retrieving a shortlist of candidate individuals from the gallery set based solely on their visual appearance, given the appearance-related content of the textual query. For a gallery of full-scene images $\mathcal{G}=\{I_{1},I_{2},\dots,I_{N}\}$ and a text $T$ describing the target person, we use appearance-related text that describes the target person’s visual traits, such as clothing, hairstyle, or accessories (e.g., A man wearing a red jacket and blue jeans is laughing), denoted as ${T}_{\text{app}}$. This ensures that the retrieval process prioritizes appearance-focused matching, excluding contextual or environmental cues at this stage.